OpenStudy (anonymous):

(5w+5)(5w-5) find product and simplify

jimthompson5910 (jim_thompson5910):

use the difference of squares rule (a+b)(a-b) = a^2 - b^2

jimthompson5910 (jim_thompson5910):

in this case a = 5w b = 5

OpenStudy (anonymous):

\[(a+b)(a-b)=a ^{2}-b ^{2}\]

OpenStudy (anonymous):

10w)w?

jimthompson5910 (jim_thompson5910):

(a+b)(a-b) = a^2 - b^2 (5w+5)(5w-5) = (5w)^2 - (5)^2 (5w+5)(5w-5) = 25w^2 - 25

OpenStudy (anonymous):

thats simplified?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

you can factor out 25 if you want, but it's pretty much expanded as much as possible

jimthompson5910 (jim_thompson5910):

I would stick to 25w^2 - 25

OpenStudy (anonymous):

could it just be w^2?

jimthompson5910 (jim_thompson5910):

no, not sure how you're getting that